How do you solve this.
Four times the sum of a number and -72 equals -16. Find the number

Answer by fractalier(6550)

Translate into algebra:
4*(x + (-72)) = -16
Now work it out...
4(x - 72) = - 16 (now divide both sides by 4...it's easiest this way)
x - 72 = -4 (now add 72)
x = 68